OpenStudy (anonymous):

How is x/(3^n) + y/(3^n) = z/(3^{n-1}) in the following equation? \[ a_{n}=\frac{6}{3^n}+\frac{9n}{3^n}=\frac{2+3n}{3^{n-1}} \]

OpenStudy (anonymous):

\[ a_{n}=\frac{6}{3^n}+\frac{9n}{3^n}=\frac{2+3n}{3^{n-1}} \]

geerky42 (geerky42):

\[ \begin{array}~ \dfrac{6}{3^n}+\dfrac{9n}{3^n} &= \dfrac{6+9n}{3^n}\\~\\&=\dfrac{3(2+3n)}{3^n} \\~\\ &=~~~~~~ \cdots \end{array} \]

geerky42 (geerky42):

Is that clear? @calyne

OpenStudy (anonymous):

oh yes thanks nvm